# Grounding Continuous Representations in Geometry: Equivariant Neural Fields

### v4

@@ Abstract
Conditional Neural Fields (CNFs) are increasingly being leveraged as continuous signal representations, by associating each data-sample with a latent variable that conditions a shared backbone Neural Field (NeF) to reconstruct the sample. However, existing CNF architectures face limitations when using this latent downstream in tasks requiring fine-grained geometric reasoning, such as classification and segmentation. We posit that this results from lack of explicit modelling of geometric information (e.g. locality in the signal or the orientation of a feature) in the latent space of CNFs. As such, we propose Equivariant Neural Fields (ENFs), a novel CNF architecture which uses a geometry-informed cross-attention to condition the NeF on a geometric variable—a latent point cloud of features—that enables an equivariant decoding from latent to field. We show that this approach induces a steerability property by which both field and latent are grounded in geometry and amenable to transformation laws: if the field transforms, the latent representation transforms accordingly—and vice versa. Crucially, this equivariance relation ensures that the latent is capable of (1) representing geometric patterns faitfhully, allowing for geometric reasoning in latent space, (2) weight-sharing over similar local patterns, allowing for efficient learning of datasets of fields. We validate these main properties in a range of tasks including classification, segmentation, forecasting and reconstruction, showing clear improvement over baselines with a geometry-free latent space. Code available here.
@@ 1 Introduction
Neural Fields (NeFs) (Xie et al., 2022) have recently gained prominence in the machine learning community as a novel representation method that models data as continuous functions. These fields, expressed as $f_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{c}$, map spatial coordinates—such as pixel locations $x\in\mathbb{R}^{2}$—to a corresponding signal, like RGB values $f_{\theta}(x)\in\mathbb{R}^{3}$, with $\theta$ representing the model’s parameters. The parameters are optimized to approximate a target signal $f$, ensuring $\forall x:f(x)\approx f_{\theta}(x)$. This capability makes NeFs effective for representing continuous spatial, spatio-temporal, and geometric data, particularly in cases where grid-based methods fall short (Dupont et al., 2022).
@@
A notable limitation of conventional CNFs, however, is a lack of explicit geometric interpretability; each field $f_{i}$ is encoded by a ”global” variable $z_{i}$, meaning for instance that any notion of locality or other explicit spatial relationships—which have proven a strong inductive bias in computer vision—is lost. For instance, in classification tasks, spatial organisation of an image’s content is crucial for understanding shape and enabling geometric reasoning (Van Quang et al., 2019); current neural fields lack such geometric inductive biases, limiting their performance in e.g. classification and generative modelling (Bauer et al., 2023; Papa et al., 2023). To this end, we propose equivariant neural fields (ENFs), a new class of NeFs that allows for the identification of continuous fields with concrete geometric representations (Fig. 1).
@@
When the goal is to utilise field representations $z_{i}$ in downstream tasks, it is crucial that these representations capture both textural/appearance information and explicit geometric information. Our approach is inspired by the idea of neural ideograms—learnable geometric representations (Vadgama et al., 2022; 2023), and addresses the pervasive issue of texture bias, which causes typical deep learning systems to overfit to textural patterns and ignore important geometric cues (Geirhos et al., 2018; Hermann et al., 2020). To address this challenge, we propose defining representations that (1) separate shape/geometry from appearance and (2) are localized in the input signal such that geometric concepts like orientation and distance are maintained from input to latent space. This necessitates that the geometric components of the representations have a meaningful structure, adhering to the same group transformation laws applicable to the fields. Geometrically, this means that distortions in the field translate to corresponding distortions in the latent space, ensuring that geometric (shape) variations are preserved and representable in latent space.
@@
This property ensures that if the field as a whole transforms, the latent will transform accordingly (Fig. 3).
@@ 3 Method
Before presenting our equivariant neural field design, we need to understand the constraints imposed by the steerability property equation 1. The key result is that for steerability, the field $f_{\theta}$ must be bi-invariant with respect to both coordinates and latents.
@@ 3.1 Equivariant Neural Fields conditioned on Geometric Attributes
To mitigate the computational overhead associated with the cross-attention between the latent points and the sampled coordinates, we propose employing a k-nearest neighbours (k-NN) approach for the cross-attention operation. Specifically, for each pixel, we first identify its k-nearest latent points and then compute cross-attention for this coordinate only over these $k$ nearest latents. This approach reduces computational cost while preserving the advantages of localized representations.
@@ 4 Experiments
For experimental details and hyperparameters we refer to Appx. C. Code for all experiments made available shortly, an implementation of ENFs may be found at this link.
@@ 4.1 Reconstruction capacity
First we evaluate our proposed ENFs on their reconstruction capabilities. •  Image data We show results for reconstruction trained with Meta-Learning on CIFAR10 Krizhevsky et al. (2009), CelebA64×${\times}64$ 64 (Liu et al., 2015) and ImageNet1K (Deng et al., 2009) using different bi-invariant attributes $\mathbf{a}_{m,i}$—resulting in equivariance to different corresponding transformation groups—in Tab. 2. We provide results for a Functa baseline, following the setup described in (Dupont et al., 2022). Notably, translational weight sharing ($\mathbf{a}^{\mathbb{R}^{2}}$) outperforms settings with no-transformation ($\mathbf{a}^{\emptyset}$) and roto-translational weight sharing ($\mathbf{a}^{\rm SE(2)}$). Moreover, it seems that locality alone is itself a useful inductive bias when moving to higher resolution, more varied images; Functa outperforms ENF $\mathbf{a}^{\emptyset}$ on CIFAR10 reconstruction, but on CelebA and ImageNet all ENF parameterizations outperform Functa. These results reinforce locality and equivariance as inductive biases in image-based continuous reconstruction tasks. •  Shape data We show shape reconstruction results (Tab. 2) on two common shape representations; voxels (3D occupancy grids) and meshes. For voxel data we take train and test splits from the 16-class ShapeNet-Part segmentation dataset (Yi et al., 2016) (which we denote ShapeNet16) and fit their corresponding voxel-based representation as occupancy function $\mathbb{R}^{3}{\rightarrow}\{0,1\}$ we use this dataset in our shape segmentation experiments detailed below. Train and test splits consists of 11.827 and 2.785 samples. For mesh data we opt instead to fit the full 55-class ShapeNetCore (v2) object dataset (Chang et al., 2015) (which we denote ShapeNet55) with 42.472 and 5.000 samples in train and test splits respectively. we follow Sitzmann et al. (2020a) and fit signed-distance functions (SDFs) $\mathbb{R}^{3}{\rightarrow}\mathbb{R}$, sampling 150k points per shape and calculating their signed distance to the mesh. Unlike Dupont et al. (2022), we found that we were unable to get reconstructions of sufficient quality with meta-learning and instead obtain latents $z$ using autodecoding on these datasets - possibly hinting at a limitation in expressivity of meta-learning for our shallow model architecture. Results show that ours as well as the baseline models struggle with accurately reconstructing the underlying shape from the SDF point clouds, we think due to the more complex optimization objective.
@@ 4.2 Downstream tasks
airplane
bag
cap
car
chair
earphone
guitar
knife
lamp
laptop
motorbike
mug
pistol
rocket
skateboard
table
@@
Where classification primarily evaluates how well the latent point clouds captures global information, we also want to evaluate ENFs performance on fine-grained tasks. As such, we evaluate on the ShapeNet part segmentation task (Yi et al., 2016). The ShapeNet-Part dataset consists of point clouds for 16 ShapeNet object-classes, each with a varying number of annotated parts for a total of 50 segmentation classes. We use the ENF $\mathbf{a}^{\mathbb{R}^{3}}$ backbone defined in the voxel reconstruction task above ($f_{\theta_{\text{recon}}}$) to obtain a latent $z^{\text{recon}}$ for a shape. Then, a second ENF ($f_{\theta_{\text{seg}}}$) is trained to map points on this shape to a one-hot encoding of their corresponding segmentation classes, i.e. $f_{\theta_{\text{seg}}}(x_{m};z^{\text{recon}})$ maps a point $x_{m}$ to its segmentation label $y_{m}$. We report mean class IoU and mean instance IoU in Tab. 8, and mIoU per class. We include dense segmentations sampled from the underlying ENF in Fig. 8. Here, somewhat surprisingly, results show ENF and Functa perform comparably in this task. Again, we think this
@@
attributable to the fact that all shapes are aligned and centered - we further investigate these results in Appx. D.2. We additionally include results for point cloud-specific architectures, and NF2Vec - a framework for self-supervised representation learning on 3D data from (non-conditional) NeFs. These results show that ENF only slightly underperforms modality-specific baselines.
Flood Map Segmentation
We evaluate and compare our model against the multi-modal U-Net proposed by Drakonakis et al. (2022). As suggested by Dupont et al. (2022) we trained Functa—unable to fit the training set with MAML—using autodecoding instead. However, we found Functa unable to generalize to test images in this complex low-data setting, collapsing to remembering the
@@
training dataset (achieving 31.5 recon PSNR and 93.7 IoU on the $256{\times}256$ train set). Results (Tab. 4, Fig. 9) show the importance of inductive biases in complex limited-data regimes. We provide results for subsampled observations to simulate missing data, showing the robustness of NeF-based methods to sparsity—ENF performs well even at $10\%$ observation rate where classical convolution-based methods fail. Moreover, we show zero-shot resolution transfer results (Tab. 5) where $f_{\theta_{\text{recon}}},f_{\theta_{\text{seg}}}$ are trained on $128{\times}128$ resolution data are deployed on $64\times 64$ and $256\times 256$ resolution data without fine-tuning, showing the resolution agnostic nature of NeF-representations.
@@ C.1 Image reconstruction, classification, segmentation, forecasting
For Cifar10 (Krizhevsky et al., 2009) reconstruction and classification we use a hidden dim of 128 with 3 heads, 25 latents of size 64, a batch size of 32 and restrict the cross-attention operator to k=4 nearest latents for each input coordinate $x$. For $\sigma_{q},\sigma_{v}$ we choose 1.0 and 3.0 respectively. We train the model for 100 epochs.
@@ C.2 Functa baseline models
For the Functa baselines (Dupont et al., 2022), we try to keep as close as possible to the setup defined by the original authors. However, we found that training deeper models (¿8 layers) lead to very unstable training. Instead, we opted to go for shallower models, up to 6 layers, with a hidden dim of 512 and latent modulation size of 512 (1024 for ImageNet1K reconstruction). We would like to note here that in all experiments, the Functa baseline has larger parameter count than the ENF models applied to each task (e.g. for CIFAR10, ENF has 522K params where Functa has 2.6M params). Although we did not explore this in-depth, it seems that the proposed ENF representation is more parameter efficient compared to the deep SIREN model defined in (Dupont et al., 2022).

### v5

@@ Abstract
Conditional Neural Fields (CNFs) are increasingly being leveraged as continuous signal representations, by associating each data-sample with a latent variable that conditions a shared backbone Neural Field (NeF) to reconstruct the sample. However, existing CNF architectures face limitations when using this latent downstream in tasks requiring fine-grained geometric reasoning, such as classification and segmentation. We posit that this results from lack of explicit modelling of geometric information (e.g. locality in the signal or the orientation of a feature) in the latent space of CNFs. As such, we propose Equivariant Neural Fields (ENFs), a novel CNF architecture which uses a geometry-informed cross-attention to condition the NeF on a geometric variable—a latent point cloud of features—that enables an equivariant decoding from latent to field. We show that this approach induces a steerability property by which both field and latent are grounded in geometry and amenable to transformation laws: if the field transforms, the latent representation transforms accordingly—and vice versa. Crucially, this equivariance relation ensures that the latent is capable of (1) representing geometric patterns faitfhully, allowing for geometric reasoning in latent space, (2) weight-sharing over similar local patterns, allowing for efficient learning of datasets of fields. We validate these main properties in a range of tasks including classification, segmentation, forecasting, reconstruction and generative modelling, showing clear improvement over baselines with a geometry-free latent space. Code attached to submission here. Code for a clean and minimal repo here.
@@ 1 Introduction
Neural Fields (NeFs) (Xie et al., 2022) have recently gained prominence in the machine learning community as a novel representation method that models data as continuous functions. These fields, expressed as $f_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{c}$, map spatial coordinates—such as pixel locations $x\in\mathbb{R}^{2}$—to a corresponding signal, like RGB values $f_{\theta}(x)\in\mathbb{R}^{3}$, with $\theta$ representing the model’s parameters. The parameters are optimized to approximate a target signal $f$, ensuring $\forall x:f(x)\approx f_{\theta}(x)$. This capability makes NeFs effective for representing continuous spatial, spatio-temporal, and geometric data, particularly in cases where grid-based methods fall short (Dupont et al., 2022). Their promise lies in serving as resolution-free representation that may be used irrespective of data resolution or discretization (Xie et al., 2022). Moreover, NeF-representations unify downstream models over different data modalities, allowing for transfer of modelling principles between data modalities that otherwise require data-specific engineering efforts (Dupont et al., 2022; Papa et al., 2023).
@@
A notable limitation of conventional CNFs, however, is a lack of explicit geometric interpretability; each field $f_{i}$ is encoded by a ”global” variable $z_{i}$, meaning for instance that any notion of locality or other explicit spatial relationships—which have proven a strong inductive bias in computer vision—is lost. Althought this global representation has inherent benefits, e.g. enabling the use of simple MLPs as downstream models and allowing for intuitive interpolation between latent signal representations, empirically it has shown limited performance in settings where samples of the dataset are not consistently globally aligned (Bauer et al., 2023). For instance, in classification tasks, spatial organisation of an image’s content is crucial for understanding shape and enabling geometric reasoning (Van Quang et al., 2019); current neural fields lack such geometric inductive biases, limiting their performance in e.g. classification and generative modelling (Bauer et al., 2023; Papa et al., 2023). To this end, we propose equivariant neural fields (ENFs), a new class of NeFs that allows for the identification of continuous fields with concrete geometric representations (Fig. 1).
@@
When the goal is to utilise field representations $z_{i}$ in downstream tasks, it is crucial that these representations capture both textural/appearance information and explicit geometric information. Our approach is inspired by the idea of neural ideograms—learnable geometric representations (Vadgama et al., 2022; 2023), and addresses the pervasive issue of texture bias, which causes typical deep learning systems to overfit to textural patterns and ignore important geometric cues (Geirhos et al., 2018; Hermann et al., 2020). To address this challenge, we propose defining representations that (1) explicitly separate aspects of geometry–specifically the pose of features–from appearance and (2) are localized in the input signal such that geometric concepts like orientation and distance are maintained from input to latent space. This necessitates that the geometric components of the representations have a meaningful structure, adhering to the same group transformation laws applicable to the fields. Geometrically, this means that distortions in the field translate to corresponding distortions in the latent space, ensuring that geometric (shape) variations are preserved and representable in latent space.
@@
This property ensures that if the field transforms, the latent transforms accordingly (Fig. 3). This property has also been shown in (Atzmon et al., 2022; Chatzipantazis et al., 2022) for learning implicit shape representations and is well known in equivariance literature more broadly (Bekkers, 2019; Cohen et al., 2019; Weiler & Cesa, 2019; Deng et al., 2021).
@@ 3 Method
Before presenting our equivariant neural field design, we need to understand the constraints imposed by the steerability property (Eq. 1). The key result–shown also by Cohen et al. (2019); Bekkers (2019) in the context of equivariant CNNs–is that for steerability, the field $f_{\theta}$ must be bi-invariant with respect to both coordinates and latents.
@@ 3.1 Equivariant Neural Fields conditioned on Geometric Attributes
To mitigate the computational overhead associated with the cross-attention between the latent points and the sampled coordinates, we propose employing a k-nearest neighbours (k-NN) approach for the cross-attention operation. Specifically, for each pixel, we first identify its k-nearest latent points and then compute cross-attention for this coordinate only over these $k$ nearest latents. This approach reduces computational cost while preserving the advantages of local representations (Appx. D.5).
@@ 4 Experiments
For experimental details and hyperparameters we refer to Appx. C.
@@ 4.1 Reconstruction capacity
First we evaluate our proposed ENFs on their reconstruction capabilities. •  Image data We show results for reconstruction trained with Meta-Learning on CIFAR10 Krizhevsky et al. (2009), CelebA64×${\times}64$ 64 (Liu et al., 2015) and ImageNet1K (Deng et al., 2009) using different bi-invariant attributes $\mathbf{a}_{m,i}$—resulting in equivariance to different corresponding transformation groups—in Tab. 2. We provide results for a Functa baseline, following the setup described in (Dupont et al., 2022). Notably, translational weight sharing ($\mathbf{a}^{\mathbb{R}^{2}}$) outperforms settings with no-transformation ($\mathbf{a}^{\emptyset}$) and roto-translational weight sharing ($\mathbf{a}^{\rm SE(2)}$). Moreover, it seems that locality alone is itself a useful inductive bias when moving to higher resolution, more varied images; Functa outperforms ENF $\mathbf{a}^{\emptyset}$ on CIFAR10 reconstruction, but on CelebA and ImageNet all ENF parameterizations outperform Functa. These results reinforce locality and equivariance as inductive biases in image-based continuous reconstruction tasks. •  Shape data We show shape reconstruction results (Tab. 2) on two common shape representations; voxels (3D occupancy grids) and meshes. For voxel data we take train and test splits from the 16-class ShapeNet-Part segmentation dataset (Yi et al., 2016) (which we denote ShapeNet16) and fit their corresponding voxel-based representation as occupancy function $\mathbb{R}^{3}{\rightarrow}\{0,1\}$ also using the obtained representations in the ShapeNet-Part segmentation experiment detailed below. For mesh data we opt instead to fit the full 55-class ShapeNetCore (v2) object dataset (Chang et al., 2015), fitting these with 150k points sampled from the signed distance function of the mesh (more details in Appx. C.3). Unlike Dupont et al. (2022), we were unable to get sufficient quality reconstructions with meta-learning and instead obtain latents $z$ using autodecoding on shape data (finding discussed in Appx. A.1.1). Results show that ours as well as the baseline models struggle with accurately reconstructing the underlying shape from the SDF point clouds, we think due to the more complex optimization objective.
@@ 4.2 Downstream tasks
Table 3: Mean class and instance IoU ($\uparrow$) on ShapeNet.
@@
PointNet 83.1 79.0
PointNet++ 84.9 82.7
DGCNN 83.6 80.9
NF2vec 81.3 76.9
Functa 82.8 74.8
ENF 82.2 75.4
@@
Where classification primarily evaluates how well the latent point clouds captures global information, we also want to evaluate ENFs performance on fine-grained tasks. As such, we evaluate on the ShapeNet part segmentation task (Yi et al., 2016). The ShapeNet-Part dataset consists of point clouds for 16 ShapeNet object-classes, each with a varying number of annotated parts for a total of 50 segmentation classes. We use the ENF $\mathbf{a}^{\mathbb{R}^{3}}$ backbone defined in the voxel reconstruction task above ($f_{\theta_{\text{recon}}}$) to obtain a latent $z^{\text{recon}}$ for a shape. Then, a second ENF ($f_{\theta_{\text{seg}}}$) is trained to map points on this shape to a one-hot encoding of their corresponding segmentation classes, i.e. $f_{\theta_{\text{seg}}}(x_{m};z^{\text{recon}})$ maps $x_{m}$ to its class label $y_{m}$. Results (Tab. 3) somewhat surprisingly show ENF and Functa perform comparably in this task (detailed results and visualizations in Appx. D.6). We again think this
@@
attributable to the fact that all shapes are aligned and centered - we further investigate these results in Appx. D.6.1. We additionally include results for point cloud-specific architectures, and NF2Vec - a framework for self-supervised representation learning on 3D data from (non-conditional) NeFs. These results show that ENF only slightly underperforms modality-specific baselines.
Flood Map Segmentation
We evaluate and compare our model against the multi-modal U-Net proposed by Drakonakis et al. (2022). As suggested by Dupont et al. (2022) we trained Functa—unable to fit the training set with MAML—using autodecoding instead. However, we found Functa unable to generalize to test images in this complex low-data setting, collapsing to remembering the training dataset (achieving 31.5 recon PSNR and 93.7 IoU on the $256{\times}256$ train set). Results (Tab. 4, Fig. 7) show the importance of inductive biases in complex limited-data regimes. We provide results for subsampled observations to simulate missing data, showing the robustness of NeF-based methods to sparsity—ENF performs well even at $10\%$ observation rate where classical convolution-based methods fail. Moreover, we show zero-shot resolution transfer results (Tab. 5) where $f_{\theta_{\text{recon}}},f_{\theta_{\text{seg}}}$ are trained on $128{\times}128$ resolution data are deployed on $64\times 64$ and $256\times 256$ resolution data without fine-tuning, showing the resolution agnostic nature of NeF-representations.
@@
Figure 9: Qualitative samples for generative modelling on CIFAR-10 and Celeba64×${\times}64$ 64.
Table 7: FID for generative modelling on CIFAR-10 and Celeba64×${\times}64$ 64.
CelebA64×${\times}64$ 64 CIFAR-10
Model FID $\downarrow$ FID $\downarrow$
GEM - 23.8
GASP 13.5 -
DPF 13.2 15.1
Functa 40.4 78.2
ENF 33.8 23.5
Image generation
Following Dupont et al. (2022); Bauer et al. (2023), we provide results for diffusion applied to a dataset of latents obtained from pretrained ENFs on CIFAR10 and Celeba64×$\times 64$ 64. As downstream diffusion model, we utilize DiT-B (Peebles & Xie, 2023), a natural choice for our set-latent (training detailed in Appx. C.4). We provide results in FID (Heusel et al., 2017) for unconditional generation in Tab. 7 and samples in Fig. 9. We provide comparison to Functa (Dupont et al., 2022), as well as other frameworks for generative modelling over fields (Du et al., 2021; Dupont et al., 2021; Zhuang et al., 2023)–notably each of these methods is trained on a generative objective and does not support self-supervised pre-training like Functa or ENF. On globally aligned CelebA64×${\times}64$ 64, both Functa and ENF produce perceptually qualitative samples, but unlike ENF, Functa is unable to generalize to CIFAR-10, where data is less homogeneous and not aligned. These results again show clear benefit of a geometrically interpretable latent space for downstream tasks, though previously proposed frameworks specific to generative modelling over fields achieve better performance than ENF. The latter points to a possible area of improvement, and future work could look into incorporating insights from these works into a generative adaptation of the ENF framework, e.g. through latent-space regularization of perceptual consistency as per Du et al. (2021).
@@ 6 Reproducibility
7 Acknowledgements
David Knigge is partially funded by Elekta Oncology Systems AB and a RVO public-private partnership grant (PPS2102). David Wessels is partially funded Ellogon.AI and a public grant of the Dutch Cancer Society (KWF) under subsidy (15059/2022-PPS2). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-9549 and EINF-10544.
@@ C.1 Image reconstruction, classification, segmentation, forecasting
For CIFAR10 (Krizhevsky et al., 2009) reconstruction and classification we use a hidden dim of 128 with 3 heads, 25 latents of size 64, a batch size of 32 and restrict the cross-attention operator to k=4 nearest latents for each input coordinate $x$. For $\sigma_{q},\sigma_{v}$ we choose 1.0 and 3.0 respectively. We train the ENF model and the classifier for 100 epochs.
@@ C.2 Functa baseline models
For the Functa baselines (Dupont et al., 2022), we try to keep as close as possible to the setup defined by the original authors. However, we found that training deeper models in the shape experiments ($\geq$ 8 layers) lead to very unstable training. Instead, for these experiments, we opted to go for shallower models, up to 6 layers. For all experiments we use a hidden dim of 512 and latent modulation size of 512 as used in (Bauer et al., 2023), except for ImageNet1K reconstruction, where we use a 1024 latent modulation. We would like to note here that in all experiments, the Functa baseline has larger parameter count than the ENF models applied to each task (e.g. for CIFAR10, ENF has 522K params where Functa has 2.6M params). Although we did not explore this in-depth, it seems that the proposed ENF representation is more parameter efficient compared to the deep SIREN model defined in (Dupont et al., 2022).
For downstream models we follow (Bauer et al., 2023) and use a 1024 hidden dim 3 layer residual MLP ($\sim$2.1M params). Like in our ENF experiments, we use the same architecture across tasks, only changing the output head to accommodate.
@@ C.3 Shape reconstruction and classification
C.4 Generative modelling on ENF latent space
For generative modelling experiments CelebA64×${\times}64$ 64 and CIFAR-10 we train a Diffusion Transformer (DiT-B) (Peebles & Xie, 2023) on ENF latents, utilizing the context vectors as input tokens and their positions as input for an RFF position embedding added to the tokens. We use the same approach for CelebA as for CIFAR10; we train an $\mathbf{a}^{\mathbb{R}^{2}}$ ENF with MAML on the image dataset, and use this model to obtain sets of ”ground truth” latents $z_{0}:=\{p_{i},\mathbf{c}_{i,0}\}_{i=1}^{N}$ for each image. We then train a DiT-B on a diffusion objective on this latent space, where the forward diffusion kernel is given by:
$z_{t}=\{(p_{i},\sqrt{\bar{\alpha}_{t}}\mathbf{c}_{i,0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{\mathbf{c}})\}_{i=1}^{N},$ (9)
with $\epsilon^{\mathbf{c}}\in N(0,1)$, i.e. we only add noise to the latent vectors, as we find adding noise to the poses leads to unstable training (something to be investigated in future work). To generate a sample, we take a random set of ”ground truth” poses from the training set, and attach a context vector $\mathbf{c}_{i,t}\in N(0,1)$ to each pose to denoise. We supervise the DiT-B with the v objective (Salimans & Ho, 2022). Like (Peebles & Xie, 2023), we use a $t_{\text{max}}{=}1000$ linear noise schedule ranging from $41e{-}$ 4 to $21e{-}$ 2, and generate samples using DDIM (Song et al., 2020) in 512 steps.
In both settings, we train the diffusion model for 200 epochs using Adam, a constant learning rate of $41e{-}$ 4, no weight-decay or dropout, and generate 50k samples to calculate FID.